ECH 4323L: Process Control Lab Laboratory 2

\underline{Objective}: To suggest hardware for automating the unit operations laboratory.

\underline{Procedure}:

An experiment from the unit operations laboratory has been assigned to your team (please see Table 1). Based on your experience with this experiment, please suggest the necessary hardware to automate this experiment. You may use the following resources to get the necessary information:

1. The website for Omega Engineering, Inc. (www.omega.com)

2. Perry’s Handbook of Chemical Engineers

3. Any reference from the library on control systems hardware

In each experiment, there are two different aspects (1) how to measure process variables and (2) how to do feedback control. Since there are six experiments and twelve teams, each experiment is assigned to two teams. One team will focus on measurement devices and the other will focus on feedback control devices.

\underline{Lab Report}: There is no written report required; however you need to make a 15 minute oral presentation on the date assigned to you in Table 1. Your presentation should cover the following aspects.

1. Brief Description of Current Experiment with a block flow diagram of the process.

2. Suggested hardware for automating the experiment.

3. Detailed description of what each piece of suggested hardware does in the experiment

4. A process flow diagram with the process control hardware in place.

5. Cost of the hardware.

The two teams that are assigned to particular experiment are allowed to collaborate; however, they have to make two separate presentations.

Team  Experiment  Focus  Presentation
Getz, Golston, Robinson  Membrane Separation Measurement March 29
Doyle, Griffin, Tart  Membrane Separation Control  March 29
Guilloty, Shaba, Stallworth  Heat Exchanger  Measurement March 29
Bostian, Lewis, T. Williams  Heat Exchanger  Control  March 29
Cornell, Daley, Davey  Extraction  Measurement March 31
Carpenter, Hemingway, Schuknecht Extraction  Control  March 31
Grossman, Kurpe, Ordetx  Absorption  Measurement March 31
Estaso, Garcia-Tunon, Maly  Absorption  Control  March 31
Hughes, Sunderhaus, Wilson  Reactor  Measurement April 5
Chin-Fook, H. Williams  Reactor  Control  April 5
George, Pace, Trotman  Distillation  Measurement April 5
Decoteau, Nwabuzor, Quarles  Distillation  Control  April 5